Pete Frank bought a computer for $4,000. Pete put down $500 and financed the balance at 10 ½% for 36 months. What is his monthly payment? (Use loan amortization table.)
Answer #1

Price of computer = $4000

down payment = $500

So, loan amount PV = $4000-500 = $3500

interest rate r = 10.5% p.a. compounded monthly

term period = 36 months

So, this is an ordinary annuity, with monthly payment of

A = PV*(r/n)/(1 - (1+r/n)^(-n*t)) = 3500*(0.105/12)/(1 - (1+0.105/12)^(-36)) = $113.76

So, monthly payment = $113.76